OpenStudy (lemanual):

16×22=◻ 212 312 28 38

OpenStudy (lemanual):

its fractions

OpenStudy (campbell_st):

simply split the 2nd number so the question is \[16 \times 20 + 16 \times 2 = \]

OpenStudy (lemanual):

1 over 6 x 2 over 2 =

OpenStudy (campbell_st):

ok... do i, multiplying fractions.. multiply the numerators multiply the denominators simplify \[\frac{1}{6} \times \frac{2}{2} = \frac{1 \times 2}{2 \times 6} = \] then simplify the answer...